OpenStudy (anonymous):

The function y=x/x^2-9 is undefined when the value of x is which of the following? A. 0 or 3 B. 3 or -3 C. 3 only D. -3 only

OpenStudy (mertsj):

What is the denominator?

OpenStudy (anonymous):

\[x^2 - 9\]

OpenStudy (mertsj):

Are there any values that a denominator cannot be?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Im confused

OpenStudy (mertsj):

Remember that dividing by 0 is meaningless so the denominator cannot be 0

OpenStudy (mertsj):

That is a general rule you should learn.

OpenStudy (anonymous):

ok

OpenStudy (mertsj):

So once we realize that it is illegal for the denominator to be 0 we need to find the illegal values of x that result in a denominator of 0

OpenStudy (anonymous):

ok `

OpenStudy (mertsj):

So now the question is : what x value or values will make x^2-9 = 0

OpenStudy (mertsj):

So I can see that something minus 9 is 0 and that something is 9.

OpenStudy (mertsj):

So if x^2 is 9, the denominator is 0.

OpenStudy (anonymous):

so a

OpenStudy (mertsj):

So if x^2 is 9, what is x?

OpenStudy (mertsj):

0^2-9 is not 0

OpenStudy (mertsj):

We could make it easier by saying: \[x^2-9=(x-3)(x+3)=0\]

OpenStudy (mertsj):

\[x-3=0 or x+3=0\]

OpenStudy (mertsj):

\[x=3 or x=-3\]

OpenStudy (anonymous):

so yues a

OpenStudy (mertsj):

So I would say that B is the correct answer.

OpenStudy (anonymous):

k